In a building with floors 1 through 7, on which floor would you be if the walk to the seventh floor is twice as long as the walk to the first floor?
Answer by richwmiller(17219)

let x = current floor
7-x is the distance from the current floor to floor 7
x-1 is the distance from the current floor to the first floor
7-x=2y
substitute
7-x=2(x-1)
7-x=2x-2
add 2 to both sides
9-x=2x
add x to both sides
9=3x
divide by 3
3=x
You are on the third floor.
check
3-1=2
7-3=4
4 is twice 2
ok